f (x, θ)-θ(1 _ θ)-1 2.7 Let X be a r.v. having th Then show that X is sufficient for θ , X-1, 2, , θ e Ω (0,1)
Answer #1

The joint probability mass function of X is

f(x_1,x_2,...,x_n|\theta ) =f(x_1|\theta )*f(x_2|\theta )*...*f(x_n|\theta )

T1,T2, ...,^n

T(1

Ti-n

Now, according to Fisher's Factorization theorem, we can write

\Rightarrow f(x_1,x_2,...,x_n|\theta ) =g(T(\underbrace{x}),\theta )*h(\underbrace{x})

where g(T(\underbrace{x}),\theta )= \theta^n(1-\theta )^{\sum_{i=1}^{n}x_i - n}*I(X_{(n)} \leq \theta ) and h(\underbrace{x}) = 1

So, TL x , θ) = X(n) will be sufficient statistic for \theta